1.   Calculate the freezing point of a solution containing 50.0 g of NaCl in 5000 mL of water. Kf for water is 1.86 C˚/m  


1
Expert's answer
2022-03-07T08:57:01-0500

Molar mass of NaCl"=58.44g\/mol"


"\Delta\>T_f=k_f×m×i"


Concentration of NaCl"=\frac{50}{5}=10g/kg"


"=\frac{10}{58.44}=0.1711"


"\Delta\>T_f=1.86×0.1711×2"

"=0.6366\u00b0C"


Freezing point "=0-0.6366"

"=-0.6366\u00b0C"